What is the value of AX after these five lines of code?

mov AX,10

shl AX,1

add AX,10

cmp AX, 10

dec ax
Answer #1

Solution:

Instruction1: mov AX,10

Explanation: Here mov is opcode and AX is called destination register value and 10 is called source register value. mov copies the source register value into destination register value, hence AX = 10

Instruction 2:shl AX,1

Explanation: shl is used to left-shift the binary value of register by some digit which is represented in second part. Here AX,1 represents the binary value of AX will be left-shifted by 1 digit so AX=10 which is 00001010 and after the execution of instruction shl AX,1 the binary value of AX will be left-shifted by 1 digit so finally it will be 00010100 which is 20 so now AX =20

Instruction 3: add AX,10

Explanation: add is opcode used to add contents of two registers and assigning final value to first register.Here AX=20 and other value=10 so AX= 20+10=30, So after 3rd instruction execution AX=30

Instruction 4: cmp AX,10

Explanation : cmp is opcode which is used to compare the two register contents which are source register at the side and destination register at the left side.

If destination register content > source register content => ZF(zero flag) = 0, CF(carry flag) =0

If destination register content < source register content => ZF(zero flag) = 0, CF(carry flag) =1

If destination register content = source register content => ZF(zero flag) = 1, CF(carry flag) =0

here AX>10 hence ZF=0,CF=0

Instruction 5:dec AX

Explanation: dec is opcode and AX is operand, dec is used to decrease the value of operand by 1

Hence AX-1 =30-1 = 29 (Final value of AX)